# and quantum information science

### v2

@@ 6.3.1 Particles on group manifolds
Explicit solutions of the Lanczos algorithm can be obtained when there are dynamical symmetries governing the time evolution on the 1D Krylov chain [21, 22, 171, 153, 172]. Since the Krylov space representation depends on both the Hamiltonian and the initial state, this emergent symmetry need not arise from the fundamental symmetries of the model. As we have seen, the Lanczos coefficients can be constructed from the survival amplitude. Within the same model, some scenarios may exhibit dynamical symmetries, while others may not. Below, we present examples in which a dynamical symmetry arises from the motion of a particle on a group manifold [153, 22].
Consider a family of fictitious Hamiltonians based on the $\mathbb{R}$) group,
$H=\alpha(L_{-1}+L_{1})+\gamma L_{0}+\delta\mathbbm{1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ [L_{0},L_{\pm 1}]=\mp L_{\pm 1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ [L_{1},L_{-1}]=2L_{0}\,.$ (83)
@@
This form also appears in other contexts such as the SYK model at large $q$. Indeed, for the growth of operators with dimension $h\sim 1/q$, we have $\gamma=0$ and $\alpha=\pi/\beta$ [153]. Quantum speed limits have also been used to show that this form leads to the fastest possible universal growth of operators [173]. More generally, (85) can be divided into three different classes depending on the relative magnitude of the parameters. The class with $\gamma<2\alpha$ where complexity grows exponentially, the class with $\gamma>2\alpha$ where complexity is periodic and the intermediate class with exactly $\gamma=2\alpha$ where the complexity grows quadratically.
@@
Evolution of the initial highest (lowest) weight state $\lvert j,-j\rangle$ labeled by spin $j$ allows us to express the dynamics using coherent states of SU(2) and discrete series representations labeled by the spin $j$. Indeed, there are $12j+$ 1 Krylov basis vectors and Lanczos coefficients that read
@@
In [171] some of these solutions were found by mapping the Schrödinger equation into the Toda system that was then solved analytically. In [172], the symmetry approach was explained by recognizing that the Schrödinger equation can be effectively written as a three-term recursion relation for orthogonal polynomials in the Liouvillian (or the Hamiltonian). Consequently, when the Lanczos coefficients have a particular dependence on $n$, which makes this equation coincide with the defining relation of known polynomial families, analytical solutions for the wave functions and complexity can be obtained. For example, the $\mathbb{R}$) family with $b_{n}$ given by (87) corresponds to Meixner orthogonal polynomials.
Figure 7: Operator growth and Krylov complexity for $\mathbb{R}$) on the Fubini-Study geometry. Operator evolution becomes a geodesic (in orange) on the hyperbolic disc. Volume (in yellow) of the region enclosed by the particle’s position is proportional to the Krylov complexity.
For the symmetry classes discussed above, Krylov complexity admits a geometrical interpretation. Namely, the dynamics on the Krylov chain can be mapped onto motion in the phase space spanned by coherent states. Using the Fubini-Study (FS) metric on the corresponding phase spaces, operator growth in the $\mathbb{R}$), SU(2), and Heisenberg-Weyl cases corresponds to particle (geodesic) motion on the hyperbolic disk, the sphere, and the complex plane, respectively. Moreover, Krylov complexity is proportional (with a factor of $\pi$) to the volume enclosed from the origin to the radius determined by the particle’s position at time $t$. An example of this for the $\mathbb{R}$) symmetry is illustrated in Fig. 7 (from [153]). This result may seem counterintuitive given the conventional intuition from Nielsen’s definition of complexity, which associates complexity with geodesic length. Nevertheless, Krylov complexities in all of the symmetry examples above exhibit this feature. Interestingly, at late times, it is the Krylov entropy (74), that scales in a manner similar to the geodesic length in the Fubini-Study geometry [153].
@@ 6.3.2 Operator growth in spin chains
Extensive numerical studies of operator growth and Krylov complexity have been conducted in [21, 152, 170, 174, 175]. In generic finite-dimensional, non-integrable systems, the growth of the Lanczos coefficients $b_{n}$ with $n$ exhibits three distinct regimes, which determine the scaling behavior of Krylov complexity. Denoting the number of degrees of freedom by $K$, [^22] In some of the references, the number of degrees of freedom is denoted by $S$ and the size of the Hilbert space by $D$. We traded those symbols for $K$ and $N$ respectively, for consistency with other parts of the review. The only exception is the section about SYK where the number of degrees of freedom is denoted $N$ for consistency with the large literature on the topic. the first regime is characterized by a linear growth of $b_{n}$’s, leading to the exponential growth of Krylov complexity. For infinite-dimensional systems, this exponential growth persists indefinitely. For finite dimensional systems, after times of order $t\geq\log(K)$ or Krylov index $n\sim K$, the Lanczos coefficients saturate at a constant value $b_{n}\sim\Lambda K$, where $\Lambda$ is a parameter related to the system’s spectral bandwidth. In this regime, Krylov complexity grows linearly with time. The above-mentioned time evolution of the Lanczos coefficients resembles a switchback effect, see Fig. 5. Indeed, a linear growth is achieved after an initial time, logarithmic in the number of degrees of freedom. Finally, at exponentially large times $t\sim e^{K}$, the Lanczos coefficients gradually decrease toward zero, leading to the saturation of Krylov complexity.
@@
Moreover, the authors of [175] examined how operator growth differs between integrable and chaotic spin-chain models. The authors studied the XXZ spin chain with integrability-breaking deformations, allowing them to interpolate between the two regimes. Noting that the behavior of Krylov complexity depends on both the Hamiltonian and the choice of initial operator, they focused on examples of initial operators such as the sum of two spin operators located at some fixed distance from the ends of the chain. In these examples, they found that the late-time plateau value is suppressed in the integrable phase, whereas in the chaotic phase, it is larger and eventually matches the prediction of random matrix theory for a system of the same dimension. The authors argued that this suppression arises from a sort of Anderson localization in the one-dimensional Krylov chain arising from disorder in the Lanczos coefficients. More details can be found in the review [3].
@@
However, at least for generic initial operators/states, the plateau value cannot serve as a signature of chaos vs. integrability, for the reasons described below. Recall from Sec. 6.1.4 that there is an analytical formula relating the density of states and the coarse-grained Lanczos spectrum. This formula applies to cases where the Hamiltonian and initial state lead to Lanczos coefficients that have a continuous large system limit, for example, if we are working with generic initial states with continuous support in the energy basis. The authors of [24] showed that for such states, the late time plateau of the spread complexity can be determined analytically in terms of the coarse-grained Lanczos spectrum. Since the relation between this spectrum and the density of states is universal, i.e., it applies to both integrable and chaotic theories which can have the same density of states, the plateau value cannot generally separate integrability and chaos, although it may do so for specially prepared initial conditions. That said, as we will see in Sec. 6.3.4, the central phenomenon driving the results of [175], i.e., order vs disorder in the Lanczos coefficients, can be distilled into a sharp conjecture regarding the covariances of the Lanczos coefficients for chaotic theories [23, 24]. These covariances affect the approach of the spread complexity to the late time plateau even in cases where the plateau value is the same between integrable and chaotic theories. We will describe this conjecture and compare its predictionsexplicitly with results for integrable and chaotic spin chains in Sec. 6.3.4.
@@ 6.3.3 SYK model
The authors of [21], originally suggested Krylov complexity as a diagnostic of chaos. They proposed that this quantity would grow exponentially in chaotic systems because of the rapid spreading of operators expected there. In particular, according to [21], if the Lanczos coeffcients grow linearly at large $n$
@@
For chaotic systems where (93) holds, [21] also proposed that the Krylov exponent upper bounds the Lyapunov exponent
@@
The first inequality was proven at infinite temperature in [21] and, under the assumption of certain analytic and smoothness properties of the Lanczos coefficients, the right inequality was proven at finite temperature in [176, 177]. The inequality $\lambda_{\text{OTOC}}\leq\frac{2\pi}{\beta}$, where $\beta=1/T$ is the inverse temperature, is the bound on chaos in quantum mechanical systems from [10] (the definition of $\lambda_{\text{OTOC}}$ is reviewed in Sec. 2).
These conjectures have been studied extensively in the SYK model, which has had important applications in the description of strange metals, the study of quantum chaos and as a model for quantum gravity, see, e.g., [178, 179, 180, 144, 165, 181, 182, 183]. The SYK Hamiltonian is constructed from $N$ Majorana fermions satisfying anti-commutation relations $\{\psi_{i},\psi_{j}\}=\delta_{ij}$ with $i,j=1,\ldots,N$:
@@
The SYK model approaches a nearly conformal fixed point at low temperatures, where the fermions acquire a conformal dimension of $1/q$ and the model becomes (nearly) maximally chaotic in the sense of the chaos bound [10]. The case $q=2$ is special – it is integrable rather than chaotic.
The SYK model admits some simplifications in the limit of infinite $N$, and even more when we take $q$ to also be large.242424 Here we mean that we first take the limit $N\rightarrow\infty$ and then we take $q\rightarrow\infty$. In the language of DSSYK, where $N$ is taken to infinity keeping the ratio $\lambda_{\text{SYK}}=2q^{2}/N$ fixed, our limit has $\lambda_{\text{SYK}}\rightarrow 0$. In this case, the survival amplitude averaged over the SYK ensemble is known analytically and one can use it to extract an averaged Lanczos sequence. At large $N$, some quantities are self averaging including the thermal two-point function [183] and many texts, including the pioneering Ref. [21], work under the assumption that the Krylov complexity is similarly self-averaging. These works extract the Lanczos sequence from the ensemble averaged survival amplitude using the moment method described in Sec. 6.1.2.252525Sometimes the survival amplitude is referred to as the auto-correlation function or thermal two-point function in this context. As we will see below in the discussion of Random Matrix Theory in Sec. 6.3.4, ensemble averages of this kind match the coarse-grained Lanczos coefficients described in Sec. 6.1.4, but do not capture higher moments such as covariances of the Lanczos spectrum. Applying the moment method to the ensemble averaged survival amplitude, the leading contribution to the Lanczos coefficients at large $q$ reads
$b_{n,\beta}^{\text{SYK}}=\begin{dcases*}2\nu\beta^{-1}\sqrt{2/q}+O(1/q)&if $n=1$\,,\\[4.30554pt] 2\nu\beta^{-1}\sqrt{n(n-1)}+O(1/q)&if $n>1$\,.\end{dcases*}$ (98)
@@
In this limit, the leading large $q$ contribution to the Krylov exponent is exactly equal to the Lyapunov exponent for the OTOC correlators [21]:
@@
A particularly interesting case is the limit of low temperatures (large $\beta$). In this case both exponents approach the chaos bound $2\pi/\beta$. The surprising equality between the Krylov and Lyaponov exponent for the OTOC suggested that the Krylov exponent, which is simpler to compute than the OTOC, would provide a shortcut for computing the Lyapunov exponent of the latter. However, at large but finite $q$, the equality fails at order $1/q$, that is [184]
@@
The perturbation is an infrared deformation of the original SYK since it has effective dimension $\tilde{q}/q<1$ around the fixed point of the original Hamiltonian. This means that as we keep flowing to the infrared, a new fixed point will appear that is dominated by the deformation. This type of deformation has been studied in, e.g., [185, 186, 187, 188, 189, 190, 191, 192, 193, 194]. In the context of chaos, we expect the infrared deformation to induce a transition between two chaotic regimes, or if $\tilde{q}=2$, between chaos and integrability along the RG flow controlled by the temperature parameter. We can try to diagnose these chaos-chaos or chaos-integrability transitions using the OTOC and Krylov exponents [184]. Fig. 8 show that while the Lyapunov exponent exhibits features indicative of such transitions, the Krylov exponent does not.
Figure 8: Lyapunov and Krylov exponents for flow SYK theories (101). Left: a flow from $q=6$ to $q=2$ at large $N$ with $s=0.2$. The Lyapunov exponent indicates that chaos develops at intermediate temperature (in the intermediate IR) but integrable behavior is approached at much lower temperatures (in the deep infrared). Right: flow between two chaotic regimes at large $N$ and large $q$ where $\tilde{q}/q=2$ is kept fixed when taking the limit and $s=10^{-3}$. We observe two regimes of near maximal chaos with a clear transition indicated by the OTOC exponent’s minima but not by the Krylov exponent. Plots adapted from [184].
@@
Additionally, the studies described above are investigating early time chaos. As we will discuss below, the spread complexity of chaotic systems grows to a peak at late times that are exponential in the entropy, and then slopes back down to a plateau. Equivalently, the peak occurs at a time linear in the dimension of the Hilbert space, which scales as the exponential of the entropy. The presence of this peak and slope depend on characteristic spectral correlations that are absent in integrable theories with Poisson spectra [22]. We will discuss this below in the context of Random Matrix Theory, but all the characteristic behavior of spread complexity that we will describe applies equally to chaotic SYK models.

### v3

@@ 6.3.1 Particles on group manifolds
Explicit solutions of the Lanczos algorithm can be obtained when there are dynamical symmetries governing the time evolution on the 1D Krylov chain Parker:2018yvk, Balasubramanian:2022tpr, Dymarsky:2019elm, Caputa:2021sib, Muck:2022xfc. Since the Krylov space representation depends on both the Hamiltonian and the initial state, this emergent symmetry need not arise from the fundamental symmetries of the model. As we have seen, the Lanczos coefficients can be constructed from the survival amplitude. Within the same model, some scenarios may exhibit dynamical symmetries, while others may not. Below, we present examples in which a dynamical symmetry arises from the motion of a particle on a group manifold Caputa:2021sib, Balasubramanian:2022tpr.
[…]
$H=\alpha(L_{-1}+L_{1})+\gamma L_{0}+\delta\mathbbm{1}\penalty 10000\ \penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \penalty 10000\ [L_{0},L_{\pm 1}]=\mp L_{\pm 1}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ ;\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ [L_{1},L_{-1}]=2L_{0}\,.$ (83)
@@
This form also appears in other contexts such as the SYK model at large $q$. Indeed, for the growth of operators with dimension $h\sim 1/q$, we have $\gamma=0$ and $\alpha=\pi/\beta$ Caputa:2021sib. Quantum speed limits have also been used to show that this form leads to the fastest possible universal growth of operators Hornedal:2022pkc. More generally, (85) can be divided into three different classes depending on the relative magnitude of the parameters. The class with $\gamma<2\alpha$ where complexity grows exponentially, the class with $\gamma>2\alpha$ where complexity is periodic and the intermediate class with exactly $\gamma=2\alpha$ where the complexity grows quadratically.
@@
Evolution of the initial highest (lowest) weight state $\lvert j,-j\rangle$ labeled by spin $j$ allows us to express the dynamics using coherent states of SU(2) and discrete series representations labeled by the spin $j$. Indeed, there are $2j+1$ Krylov basis vectors and Lanczos coefficients that read
@@
[…]
@@ 6.3.2 Operator growth in spin chains
Extensive numerical studies of operator growth and Krylov complexity have been conducted in Parker:2018yvk, Barbon:2019wsy, Rabinovici:2020ryf, Rabinovici:2021qqt, Rabinovici:2022beu. In generic finite-dimensional, non-integrable systems, the growth of the Lanczos coefficients $b_{n}$ with $n$ exhibits three distinct regimes, which determine the scaling behavior of Krylov complexity. Denoting the number of degrees of freedom by KK, [^25] In some of the references, the number of degrees of freedom is denoted by $S$ and the size of the Hilbert space by $D$. We traded those symbols for $K$ and $N$ respectively, for consistency with other parts of the review. The only exception is the section about SYK where the number of degrees of freedom is denoted $N$ for consistency with the large literature on the topic. the first regime is characterized by a linear growth of $b_{n}$’s, leading to the exponential growth of Krylov complexity. For infinite-dimensional systems, this exponential growth persists indefinitely. For finite dimensional systems, after times of order $t\geq\log(K)$ or Krylov index $n\sim K$, the Lanczos coefficients saturate at a constant value $b_{n}\sim\Lambda K$, where $\Lambda$ is a parameter related to the system’s spectral bandwidth. In this regime, Krylov complexity grows linearly with time. The above-mentioned time evolution of the Lanczos coefficients resembles a switchback effect, see Fig. 5. Indeed, a linear growth is achieved after an initial time, logarithmic in the number of degrees of freedom. Finally, at exponentially large times $t\sim e^{K}$, the Lanczos coefficients gradually decrease toward zero, leading to the saturation of Krylov complexity.
@@
Moreover, the authors of Rabinovici:2022beu examined how operator growth differs between integrable and chaotic spin-chain models. The authors studied the XXZ spin chain with integrability-breaking deformations, allowing them to interpolate between the two regimes. Noting that the behavior of Krylov complexity depends on both the Hamiltonian and the choice of initial operator, they focused on examples of initial operators such as the sum of two spin operators located at some fixed distance from the ends of the chain. In these examples, they found that the late-time plateau value is suppressed in the integrable phase, whereas in the chaotic phase, it is larger and eventually matches the prediction of random matrix theory for a system of the same dimension. The authors argued that this suppression arises from a sort of Anderson localization in the one-dimensional Krylov chain arising from disorder in the Lanczos coefficients. More details can be found in the review Nandy:2024htc.
@@
However, at least for generic initial operators/states, the plateau value cannot serve as a signature of chaos vs. integrability, for the reasons described below. Recall from Sec. 6.1.4 that there is an analytical formula relating the density of states and the coarse-grained Lanczos spectrum. This formula applies to cases where the Hamiltonian and initial state lead to Lanczos coefficients that have a continuous large system limit, for example, if we are working with generic initial states with continuous support in the energy basis. The authors of Balasubramanian:2023kwd showed that for such states, the late time plateau of the spread complexity can be determined analytically in terms of the coarse-grained Lanczos spectrum. Since the relation between this spectrum and the density of states is universal, i.e., it applies to both integrable and chaotic theories which can have the same density of states, the plateau value cannot generally separate integrability and chaos, although it may do so for specially prepared initial conditions. That said, as we will see in Sec. 6.3.4, the central phenomenon driving the results of Rabinovici:2022beu, i.e., order vs disorder in the Lanczos coefficients, can be distilled into a sharp conjecture regarding the covariances of the Lanczos coefficients for chaotic theories Balasubramanian:2022dnj, Balasubramanian:2023kwd. These covariances affect the approach of the spread complexity to the late time plateau even in cases where the plateau value is the same between integrable and chaotic theories. We will describe this conjecture and compare its predictionsexplicitly with results for integrable and chaotic spin chains in Sec. 6.3.4.
@@ 6.3.3 SYK model
The authors of Parker:2018yvk, originally suggested Krylov complexity as a diagnostic of chaos. They proposed that this quantity would grow exponentially in chaotic systems because of the rapid spreading of operators expected there. In particular, according to Parker:2018yvk, if the Lanczos coeffcients grow linearly at large $n$
@@
For chaotic systems where (93) holds, Parker:2018yvk also proposed that the Krylov exponent upper bounds the Lyapunov exponent
@@
The first inequality was proven at infinite temperature in Parker:2018yvk and, under the assumption of certain analytic and smoothness properties of the Lanczos coefficients, the right inequality was proven at finite temperature in Avdoshkin:2019trj, Gu:2021xaj. The inequality $\lambda_{\text{OTOC}}\leq\frac{2\pi}{\beta}$, where $\beta=1/T$ is the inverse temperature, is the bound on chaos in quantum mechanical systems from Maldacena:2015waa (the definition of $\lambda_{\text{OTOC}}$ is reviewed in Sec. 2).
These conjectures have been studied extensively in the SYK model, which has had important applications in the description of strange metals, the study of quantum chaos and as a model for quantum gravity, see, e.g., Sachdev:1992fk, Kitaev:2015, Maldacena:2016hyu, Cotler:2016fpe, Roberts:2018mnp, Sarosi:2017ykf, Rosenhaus:2018dtp, Chowdhury:2021qpy. The SYK Hamiltonian is constructed from $N$ Majorana fermions satisfying anti-commutation relations $\{\psi_{i},\psi_{j}\}=\delta_{ij}$ with $i,j=1,\ldots,N$:
@@
The SYK model approaches a nearly conformal fixed point at low temperatures, where the fermions acquire a conformal dimension of $1/q$ and the model becomes (nearly) maximally chaotic in the sense of the chaos bound Maldacena:2015waa. The case $q=2$ is special – it is integrable rather than chaotic.
The SYK model admits some simplifications in the limit of infinite $N$, and even more when we take $q$ to also be large. [^27] Here we mean that we first take the limit $N\rightarrow\infty$ and then we take $q\rightarrow\infty$. In the language of DSSYK, where $N$ is taken to infinity keeping the ratio $\lambda_{\text{SYK}}=2q^{2}/N$ fixed, our limit has $\lambda_{\text{SYK}}\rightarrow 0$. In this case, the survival amplitude averaged over the SYK ensemble is known analytically and one can use it to extract an averaged Lanczos sequence. At large $N$, some quantities are self averaging including the thermal two-point function Chowdhury:2021qpy and many texts, including the pioneering Ref. Parker:2018yvk, work under the assumption that the Krylov complexity is similarly self-averaging. These works extract the Lanczos sequence from the ensemble averaged survival amplitude using the moment method described in Sec. 6.1.2.282828Sometimes the survival amplitude is referred to as the auto-correlation function or thermal two-point function in this context. As we will see below in the discussion of Random Matrix Theory in Sec. 6.3.4, ensemble averages of this kind match the coarse-grained Lanczos coefficients described in Sec. 6.1.4, but do not capture higher moments such as covariances of the Lanczos spectrum. Applying the moment method to the ensemble averaged survival amplitude, the leading contribution to the Lanczos coefficients at large $q$ reads
[…]
@@
In this limit, the leading large $q$ contribution to the Krylov exponent is exactly equal to the Lyapunov exponent for the OTOC correlators Parker:2018yvk:
@@
A particularly interesting case is the limit of low temperatures (large $\beta$). In this case both exponents approach the chaos bound $2\pi/\beta$. The surprising equality between the Krylov and Lyaponov exponent for the OTOC suggested that the Krylov exponent, which is simpler to compute than the OTOC, would provide a shortcut for computing the Lyapunov exponent of the latter. However, at large but finite $q$, the equality fails at order $1/q$, that is Chapman:2024pdw
@@
The perturbation is an infrared deformation of the original SYK since it has effective dimension $\tilde{q}/q<1$ around the fixed point of the original Hamiltonian. This means that as we keep flowing to the infrared, a new fixed point will appear that is dominated by the deformation. This type of deformation has been studied in, e.g., Garcia-Garcia:2017bkg, Kim:2020mho, Garcia-Garcia:2020dzm, Lunkin:2020tbq, Nandy:2022hcm, Menzler:2024ifs, Jiang:2019pam, Anninos:2020cwo, Anninos:2022qgy, Louw:2023lpq. In the context of chaos, we expect the infrared deformation to induce a transition between two chaotic regimes, or if $\tilde{q}=2$, between chaos and integrability along the RG flow controlled by the temperature parameter. We can try to diagnose these chaos-chaos or chaos-integrability transitions using the OTOC and Krylov exponents Chapman:2024pdw. Fig. 8 show that while the Lyapunov exponent exhibits features indicative of such transitions, the Krylov exponent does not.
Figure 8: Lyapunov and Krylov exponents for flow SYK theories (101). Left: a flow from $q=6$ to $q=2$ at large NN with $s=0.2$. The Lyapunov exponent indicates that chaos develops at intermediate temperature (in the intermediate IR) but integrable behavior is approached at much lower temperatures (in the deep infrared). Right: flow between two chaotic regimes at large NN and large $q$ where $\tilde{q}/q=2$ is kept fixed when taking the limit and $s=10^{-3}$. We observe two regimes of near maximal chaos with a clear transition indicated by the OTOC exponent’s minima but not by the Krylov exponent. Plots adapted from Chapman:2024pdw.
@@
Additionally, the studies described above are investigating early time chaos. As we will discuss below, the spread complexity of chaotic systems grows to a peak at late times that are exponential in the entropy, and then slopes back down to a plateau. Equivalently, the peak occurs at a time linear in the dimension of the Hilbert space, which scales as the exponential of the entropy. The presence of this peak and slope depend on characteristic spectral correlations that are absent in integrable theories with Poisson spectra Balasubramanian:2022tpr. We will discuss this below in the context of Random Matrix Theory, but all the characteristic behavior of spread complexity that we will describe applies equally to chaotic SYK models.